A swimming pool is rectangular shape of length 12 m and width 5.5 m. It has a sloping bottom and is 1 m deep at one end and 3.6 m deep at the other end. The water from a full cylindrical reservoir 3.6 m in diameter and 10 m deep is emptied to the pool. Find the depth of water at the deep end.


1
Expert's answer
2021-05-30T21:42:01-0400

Ans:-

The water from a full cylindrical reservoir 3.6 m in diameter and 10 m deep is emptied to the pool.

Volume of water in Cylinder

"V=\pi r^2 h = \ 3.14\times 1.8\times 10 =56.52\ m^3"

The cross section of the swimming pool is a trapezium with parallel sides "1 \ m \ and \ 3.6 \ m" (depth of both ends) and"\ 12\ m" (length of pool) as the ⊥ r distance between ∥ sides.

The width of pool "5.5" m can be the length of cross section

Capacity of Swimming pool will be

"C=\dfrac{1}{2}[(1+3.6)\times 12]\times5.5=151.8\ m^3"

Volume of Empty swimming pool "151.8-56.52=95.28\ m^3"


Hence The height of depth side from the top will be

"95.28=\dfrac{1}{2}[(1+x)\times 12 ]\times5.5 \Rightarrow x=1.887\ m"

Hence depth of water at the bottom end will be "3.6-1.887 \ =\ 1.713\ m"